Find fx(x,y), fy(x,y), fx(1,3), and fy(-2,4) for the given function. If

𝑧 = 𝑓(π‘₯, 𝑦) = 3π‘₯^3𝑦^2 βˆ’ π‘₯^2𝑦^3 + 4π‘₯ + 9


1
Expert's answer
2021-06-10T02:55:50-0400
"\ud835\udc67 = \ud835\udc53(\ud835\udc65, \ud835\udc66) = 3\ud835\udc65^3\ud835\udc66^2 \u2212 \ud835\udc65^2\ud835\udc66^3 + 4\ud835\udc65 + 9"


"f_x(x, y)=9x^2y^2-2xy^3+4"

"f_y(x, y)=6x^3y-6x^2y^2"


"f_x(1, 3)=9(1)^2(3)^2-2(1)(3)^3+4=29"

"f_y(-2, 4)=6(-2)^3(4)-6(-2)^2(4)^2=-576"